OpenStudy (anonymous):

Factor the trinomial below. x2 - 16x + 28

OpenStudy (callisto):

x^2 - 16x + 28 = x^2 - 2x -14x + 28 Then take out the common factor = x(x- 2) -14(x - 2) Take out the common factor again, what would you get?

OpenStudy (anonymous):

(x - 2)(x - 14)

OpenStudy (callisto):

Yes :)